three goats eat 6 tin cans in 1 hour . how long should it take six goats to eat
3 tin cans ?

Answer by nerdybill(7384)

three goats eat 6 tin cans in 1 hour . how long should it take six goats to eat
3 tin cans ?
.
First determine what ONE goat can do:
6 tins per 1 hour per 3 goats
For 1 goat 2 tins per 1 hour
.
Let t = time it takes
then
6(2/1)t = 3
(12)t = 3
t = 3/12
t = 1/4 hour (or 15 minutes)